OpenStudy (moonlitfate):

Find the derivative of ln(x^3+3x)^3

OpenStudy (anonymous):

depends on what is being cubed

OpenStudy (anonymous):

if it is \[\ln((x^3+3x)^3)\] then the first step would be to rewrite it as \[3\ln(x^3+3x)\]

OpenStudy (anonymous):

f(x)=[3ln(x^3+3x)*(3x^2+x)]/(x^3+3x)

OpenStudy (anonymous):

then it will be real easy, since \[\frac{d}{dx}[\ln(f(x))]=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

if the whole thing is being cubed, then the answer will be different

OpenStudy (anonymous):

sorry, it should be f(x)=[3ln(x^3+3x)^2*(3x^2+x)]/(x^3+3x)

OpenStudy (moonlitfate):

\[\ln(x^3+3x)^3\]

OpenStudy (moonlitfate):

@satellite73 So, would it be (3(x^3+3x)^2+3x^2+3)/ln(x^3+3x)^3 Too many threes. @w@

OpenStudy (anonymous):

it really depends on what is being raised to the power of 3, the input of the log, or the whole thing

OpenStudy (anonymous):

that is, is it \[\left(\ln(x^3+3x)\right)^3\] or is it \[\ln((x^3+3x)^3)\]

OpenStudy (anonymous):

if it is the first one, you get \[3\ln(x^3+3x)^2\times \frac{3x^2+3}{x^3+3x}\]

OpenStudy (moonlitfate):

It's the second thing-- the ln isn't cubed.

OpenStudy (anonymous):

if it is the second one, you get \[3\times \frac{3x^2+3}{x^3+3x}\]

OpenStudy (moonlitfate):

I tried to work it out and I got:\[\frac{ 9x^2+9 }{ (x^3+3x)^3 }\]

OpenStudy (anonymous):

the denominator is wrong

OpenStudy (moonlitfate):

Oh, it's just the inside function on the bottom?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the three comes out front as a multiplier by the property of the log that says \[\ln(x^n)=n\ln(x)\]